Circular No. 7343

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMETS C/1998 G3 AND C/1999 Y2 K. Černis, Vilnius, Lithuania, reports his discovery of an apparent

comet (discovery position below) at mag ∼ 5 on SOHO images taken on 1999 Dec. 28 that were posted on the SOHO website. D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center, reports that the comet was visible on both LASCO C2 and C3 images and that no tail was detected. Astrometric measurments by Biesecker and D. Hammer (University of Maryland), reduced by Marsden, appear on MPEC 2000-A36, together with parabolic orbital elements (q = 0.048 AU, i = 111^{o}.4), showing that the comet is not a Kreutz sungrazer. Magnitude reductions by Biesecker and Hammer show that the comet faded from mag 6.1 to 6.8 during Dec. 28.58–28.79, and thence from view while still in the C3 field.

1999 UT  α_{2000}  δ_{2000}
Dec. 28.279  18^{h}28.^{m}3  −26^{o}10′

New astrometric measurements and orbital elements for C/1998 G3 (cf. IAUC 6952) appear on MPEC 2000-A37. This comet is also not a Kreutz sungrazer.

QY PERSEI T. Kato, Kyoto University, reports on behalf of the VSNET QY Per

collaboration team: “Unfiltered CCD photometry of QY Per has been undertaken since 1999 Dec. 29.356 UT by numerous observers around the world (T. Kato, M. Uemura, K. Matsumoto, G. Masi, A. Cassetti, L. Cook, L. T. Jensen, B. Martin, T. Vanmunster, and D. Buczynski). Analysis of the combined light curve (covering 8 days) revealed the clear existence of superhumps, with an average period of 0.07861 day, including the prominent growth of superhumps starting around Dec. 29.7. The fairly long outburst interval and the long superhump period suggest that QY Per does indeed belong (cf. IAUC 7340) to a relatively rare class of long-period SU UMa-type dwarf novae. The light curve can be seen at

http://www.kusastro.kyoto-u.ac.jp/vsnet/DNe/qyper.html.”

V1494 AQUILAE Visual magnitude estimates: 1999 Dec. 26.699 UT, 7.7 (R. J. Bouma,

Groningen, The Netherlands); 29.727, 7.6 (Bouma); 2000 Jan. 3.714, 8.1 (S. Baroni, Milan, Italy); 5.702, 8.0 (K. Hornoch, Lelekovice, Czech Republic).

2000 January 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green